HETAV PATEL

Grade 11,

A particle is moving eastwards with a velocity of 5 m/s.If in 10s the velocity changes by 5 m/s northwards,what is the average acceleration in this time?

A particle is moving eastwards with a velocity of 5 m/s.If in 10s the velocity changes by 5 m/s northwards,what is the average acceleration in this time?

Grade:12th pass

1 Answers

Khimraj
5 years ago
initial velocity = 5\hat{i}
final velocity = 5\hat{j}
\Delta v = 5\hat{j} – 5\hat{i}
magnitude of change in velocity  |\Delta v| = \sqrt{5^{2} + 5^{2}} = 5\sqrt{2}
So average acceleration =  |\Delta v|/t = 5\sqrt{2}/10 = 1/\sqrt{2} m/s2
Hope it clears.